# Policy Contrastive Decoding for Robotic Foundation Models

### v2

@@ 3.2 Policy Contrastive Decoding (PCD)
where $C^{\prime}$ is a normalization constant and $\mathcal{K}(\cdot)$ indicates a Gaussian kernel $\mathcal{K}(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}$, $b$ is the bandwidth parameter controlling the smoothness of the distribution, and $\{a_{t}(j)\}_{j=1}^{N}$ are the $N$ candidate actions for the $t$-th dimension of the action $\boldsymbol{a}_{i}$, generated from the diffusion process. Similarly, we can derive an action probability distribution for $a_{t}$ conditioned on object-masked $\hat{\boldsymbol{o}}_{i}$, i.e., $\pi_{\theta}(a_{t}|l,\hat{\boldsymbol{o}}_{i})$. With the independence assumption, from $\pi_{\theta}(a_{t}|l,\boldsymbol{o}_{i})$ and $\pi_{\theta}(a_{t}|l,\hat{\boldsymbol{o}}_{i})$ we can compute the joint action distributions, which are then used to compute the contrastive action probability $\pi_{\theta}^{*}(\boldsymbol{a}_{i}|l,\boldsymbol{o}_{i})$ using Eq. (2). In this way, our PCD is compatible with diffusion-based robot policies.
@@
Inference with PCD. As a training-free method, our proposed PCD can be employed as a plugin to improve the inference stage of different types of robot policies (i.e., autoregressive policies and diffusion-based policies). Specifically, for a pre-trained robot policy $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$, given the current observation ${\boldsymbol{o}}_{i}$, we can obtain the object-masked $\boldsymbol{o}_{i}$ (i.e., $\hat{\boldsymbol{o}}_{i}$) leveraging the devised Track2Mask strategy. Based on $\boldsymbol{o}_{i}$, $\hat{\boldsymbol{o}}_{i}$ and a language command $\ell$, the autoregressive policy directly computes contrastive action probability distributions for deriving $\boldsymbol{a}_{i}$ (w/ Eq. (2)), whereas the diffusion-based policy requires performing the KDE-PM process on its output predictions to achieve the same goal. Pseudocode for our PCD method is provided in Algorithm 1.
@@
1:A pre-trained robot policy: $\pi_{\theta}(\boldsymbol{a}_{i}\mid\boldsymbol{o}_{i},\ell)$; initial observation $\boldsymbol{o}_{0}$; language command $\ell$; maximum time step $S$.
2:$s\leftarrow 0$
@@
4:     Obtain object-masked $\boldsymbol{o}_{s}$: $\hat{\boldsymbol{o}}_{s}$triangleright w/ Track2Mask
5:     Obtain $\pi_{\theta}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ and $\pi_{\theta}(\boldsymbol{a}_{s}|l,\hat{\boldsymbol{o}}_{s})$triangleright w/ KDE-PM for diffusion-based policies
6:     Compute $\pi_{\theta}^{*}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ using Eq. (2)
7:     Sample $\,\boldsymbol{a}_{s}$ from $\pi_{\theta}^{*}(\boldsymbol{a}_{s}|l,\boldsymbol{o}_{s})$ and execute $\boldsymbol{a}_{s}$
8:     $\boldsymbol{o}_{s+1}\leftarrow\text{New observation}$triangleright Obtain a new observation
@@ 4 Experiments
In this section, we conduct extensive experiments in both simulation and real-world environments, on top of three state-of-the-art open-source robot policies to answer the following questions: {mdframed} [skipabove=4pt, innertopmargin=4pt, innerbottommargin=4pt] 1) Can PCD improve robot policies in both simulation and real-world environments?
@@
3) What kinds of spurious correlations can our PCD tackle? We answer the first question in Sections 4.1 and 4.2, the second question in Section 4.3, and the third question in Section 4.4. Detailed descriptions of the baseline robot policies and the evaluation tasks used in simulation and real-world environments are presented in Appendix A.3.
@@ 4.1 Simulation Experiments
We perform simulation experiments in SIMPLER [24], a real-to-sim evaluation environment designed specifically for real-robot policies. SIMPLER accurately reflects real-world performance, enabling reliable assessment of robotic policies. We evaluate 9 tasks across two robot platforms: 5 tasks using the Google Robot and 4 tasks using the WidowX arm.
Experimental Setup. We conduct experiments on top of three diverse robot policies, including the autoregressive policy OpenVLA [20], and the diffusion-based policies Octo (base) [32] and $\pi_{0}$ [3]. We treat these policies as black-boxes, applying PCD solely based on the output action probability distributions for OpenVLA and the output actions for Octo and $\pi_{0}$. For our PCD, we annotate the target object in the initial observation using three manners: artificial Point and Box prompts, and automatic detection results from Grounding DINO [28]. The number of the sampled noise vectors $N$ in KDE-PM is set to 24 for Octo and $\pi_{0}$. We perform ablation studies on the hyperparameter $\alpha$ in Eq. (2), as well as on the object detection strategies and object inpainting paradigms of the devised Track2Mask module in Section 4.3.
Results and Analysis. Table 1 reports the performance of three robot policies (i.e., OpenVLA [20], Octo [32] and $\pi_{0}$ [3]) integrated w/ or w/o our PCD across a total of 9 tasks from Google Robot and WidowX. From the results in the table, we have the following key observations. 1) Our PCD consistently improves the three baseline policies over the 9 tasks from both the Google Robot and WidowX robotic platforms. On average, PCD boosts the success rates of the autoregressive policy OpenVLA, and the diffusion-based policies Octo and $\pi_{0}$ by up to 41%, 31% and 8%, respectively. 2) Our PCD exhibits stable and outstanding performance across the three object annotation strategies, including artificial Point and Box prompts and automatic object detection with Grounding DINO, underscoring its effectiveness and adaptability. In particular, on the Octo and OpenVLA baseline policies, leveraging the off-the-shelf model Grounding DINO demonstrates better performance than using human Point and Box prompts, while showing comparable results on $\pi_{0}$, highlighting the practical applicability of the PCD method. 3) Compared to OpenVLA and Octo, the baseline model $\pi_{0}$ demonstrates significantly superior performance across all 9 tasks. For instance, OpenVLA and Octo achieve near-zero success rates on both“Apple Drawer” and “Stack Cube” tasks whereas $\pi_{0}$ achieve success rates of 11% and 67% on these two tasks, respectively. Despite this, from the average results over the 9 tasks, our PCD still enhances the strong baseline $\pi_{0}$ by 8%. In summary, the obtained results in the table demonstrate that our PCD approach can be used as a plugin to improve different types of robot policies.
@@ 4.2 Real-world Experiments
Table 1: SIMPLER Performance—Success rates of the autoaggressive policy OpenVLA [20] and the diffusion-based policies Octo [32] and $\pi_{0}$ [3] integrated w/ or w/o PCD on 9 tasks from Google Robot and WidowX. Task-specific objects in the initial observation are annotated by artificial Point and Box prompts or the automatic detection results of Grounding DINO [28]. As a plug-and-play approach, PCD consistently enhances the three policies by large margins over the 9 tasks.
OpenVLA[20] Octo [32] $\pi_{0}$ [3]
Tasks Base +PCD Base +PCD Base +PCD
Point Box GDINO Point Box GDINO Point Box GDINO
Average 0.17 0.24 0.21 0.24 0.16 0.21 0.20 0.19 0.65 0.70 0.69 0.69
+41% +24% +41% +31% +25% +19% +8% +6% +6%
Google Robot Close Drawer 0.41 0.68 0.66 0.77 0.33 0.28 0.38 0.16 0.75 0.72 0.76 0.71
Move Near 0.58 0.62 0.52 0.53 0.07 0.13 0.12 0.13 0.73 0.74 0.71 0.75
Open Drawer 0.20 0.29 0.29 0.26 0.00 0.01 0.00 0.02 0.39 0.49 0.50 0.52
Pick Coke Can 0.25 0.40 0.42 0.42 0.28 0.54 0.49 0.52 0.84 0.88 0.87 0.87
Apple Drawer 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.29 0.24 0.22
WidX. Carrot Plate 0.00 0.08 0.00 0.04 0.14 0.28 0.22 0.19 0.59 0.67 0.62 0.62
Eggplant Basket 0.01 0.02 0.02 0.11 0.48 0.41 0.38 0.44 0.93 0.86 0.87 0.88
Spoon Towel 0.00 0.02 0.00 0.03 0.10 0.15 0.14 0.15 0.83 0.89 0.88 0.86
Stack Cube 0.04 0.01 0.00 0.00 0.00 0.05 0.06 0.07 0.67 0.77 0.78 0.75
@@
Figure 3: Real-world Performance—Success rates and time costs of the state-of-the-art robot policy $\pi_{0}$ [3] integrated w/ or w/o PCD on 6 real-world tasks. The target objects in the initial observation are automatically annotated by Grounding DINO [28]. PCD delivers a remarkable 108% performance improvement on the baseline policy, though it incurs a 24% increase in time cost.
@@
Experimental Setup. The simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA. Therefore, we employ $\pi_{0}$ as the baseline policy to validate the effectiveness of our PCD approach, as shown in Fig. 3. Considering the generalization limitations of existing robot policies in real-world environments, we first finetune $\pi_{0}$ on downstream real-world tasks before evaluating PCD’s performance. We design 6 manipulation tasks that focus on evaluating the policy’s performance along multiple dimensions: spatial reasoning and interacting with various objects and scenes. The $\pi_{0}$ policy is jointly fine-tuned on the 6 real-world tasks, with each task consisting of 10 training demonstrations. During inference, we conduct 20 trials for each real-world task and randomize the configurations and orientations of task-specific objects for each trial. We use an AGILEX PIPER 6DOF robot arm in our experiments. For our PCD, we leverage the off-the-shelf model Grounding DINO to detect task-specific objects in the initial observation, and use LaMa [39] to inpaint those objects segmented by the SAM2 [21] model. All other hyperparameters are kept consistent with the settings applied in the simulation experiments.
@@
Results and Analysis. Fig. 3 reports the success rates as well as time costs of the state-of-the-art robot policy $\pi_{0}$ [3] integrated w/ or w/o our PCD on 6 real-world tasks. For each task, we record the time cost (in seconds) from the start of the inference process to the successful completion of the task. From the obtained results in the figure, we have several key observations. 1) PCD consistently improves the success rates of the strong baseline policy across six real-world tasks, achieving an average enhancement of 108%. 2) While PCD significantly boosts success rates, it also increases the time cost by 24% on average compared to the baseline policy. However, considering the notable improvement in task completion success rates, this trade-off is acceptable in a wide range of real-world robotic applications. 3) PCD demonstrates consistent performance improvements across tasks of varying complexity. For example, in the challenging "Stack Cube" task, PCD improves the success rate from 0.05 to 0.10, showing its strong adaptability and robustness across diverse scenarios. Additionally, the backgrounds of these six real-world tasks are far more complex than those of the nine simulation tasks, containing numerous distracting elements such as plastic bottles, baskets, and trash cans. Nevertheless, our PCD shows remarkable advantages. It is worth mentioning that we use consistent PCD hyperparameters across tasks in both the simulation and real-world environments, without carefully tuning task-specific PCD hyperparameters (e.g., $\alpha$) for each individual task, leaving room for potential performance improvements through task-specific parameter tuning.
@@ 4.3 Ablation Studies
Effect of Off-the-shelf Object Detection Models. Beyond leveraging artificial Point and Box prompts, the devised Track2Mask module can employ off-the-shelf open vocabulary object detection models to annotate task-specific objects in the initial observation of the trajectory during inference. Fig. 4 (b) presents an ablation study on various automatic object detection models, including the Grounding DINO [28] model discussed in the main paper, as well as the open vocabulary object detection model YOLO World [8] and semantic segmentation model SED [48]. The output of these models is fed into the SAM2 [21] model for tracking and segmenting the detected objects across subsequent observations in the trajectory. As can be seen, our PCD consistently improves the baseline policies when integrated with each of the three off-the-shelf models, showcasing its effectiveness and stability. Generally, the Grounding DINO model yields the best results among the three models.
Effect of Object Inpainting Strategies. Once the objects specified in the language instruction of the target task are segmented by the SAM2 model, they are inpainted to create object-masked observations. In this experiment, we investigate the effect of different object inpainting strategies on PCD’s performance. Fig. 4 (c) shows the results of PCD integrated with three inpainting strategies: Telea [41] and Navier-Stokes [2] and LaMa [39]. As indicated in the figure, our PCD approach consistently improves the baseline policies across all strategies, showcasing its effectiveness and stability. Furthermore, PCD demonstrates low sensitivity to variations in the three object inpainting strategies, further validating its applicability and robustness. Notably, LaMa delivers the best results across all three baseline policies.
@@ 4.4 Robustness to Various Kinds of Spurious Correlations
The central concept of our proposed PCD approach in this work is to overcome the adverse effects of spurious features on the decision-making of robot policies. So, what types of spurious correlations can PCD effectively address? To answer the question, we evaluate the robustness of the OpenVLA [20] and $\pi_{0}$ [3] policies to task scene variations in both simulation and real-world environments. As illustrated in Fig. 5, we report the performance of the two baseline policies integrated w/ or w/o our proposed PCD approach in different testing scenarios with unseen spatial relationships, brightness, distractors and table textures within visual observations. The results shown in the figure lead to the following observations. 1) Consistent with the previous experimental results, our PCD approach significantly improves the performance of the baseline policies on both simulation and real-world tasks (Orange line vs. Blue line). 2) Both baseline policies experience substantial drops in performance when testing scenarios are modified. For example, altering the brightness of visual observations results in 48% and 75% decreases in success rates for the two policies, respectively. 3) PCD consistently alleviates the performance degradation of the baseline policies across all unseen testing scenarios. Remarkably, after integrating PCD, the two baseline policies even performs better in 4/10 of unseen scenarios than in the original training scenarios. The results in the figure reveal that our proposed PCD approach can serve as a plugin to mitigate the adverse effects of various types of spurious correlations on the generalization of robot policies.
[…]

### v3

@@ 3.2 Policy Contrastive Decoding (PCD)
where $C^{\prime}$ is a normalization constant and $\mathcal{K}(\cdot)$ indicates a Gaussian kernel $\mathcal{K}(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}$, $b$ is the bandwidth parameter controlling the smoothness of the distribution, and $\{a_{t}(j)\}_{j=1}^{N}$ are the NN candidate actions for the $t$-th dimension of the action $\bm{a}_{i}$, generated from the diffusion process. Similarly, we can derive an action probability distribution for $a_{t}$ conditioned on object-masked $\hat{\bm{o}}_{i}$, i.e., $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$. With the independence assumption, from $\pi_{\theta}(a_{t}|l,\bm{o}_{i})$ and $\pi_{\theta}(a_{t}|l,\hat{\bm{o}}_{i})$ we can compute the joint action distributions, which are then used to compute the contrastive action probability $\pi_{\theta}^{*}(\bm{a}_{i}|l,\bm{o}_{i})$ using Eq. (2). In this way, our PCD is compatible with diffusion-based robot policies.
@@
Inference with PCD. As a training-free method, our proposed PCD can be employed as a plugin to improve the inference stage of different types of robot policies (i.e., autoregressive policies and diffusion-based policies). Specifically, for a pre-trained robot policy $\pi_{\theta}(\bm{a}_{i}\mid\bm{o}_{i},\ell)$, given the current observation ${\bm{o}}_{i}$, we can obtain the object-masked $\bm{o}_{i}$ (i.e., $\hat{\bm{o}}_{i}$) leveraging the devised Track2Mask strategy. Based on $\bm{o}_{i}$, $\hat{\bm{o}}_{i}$ and a language command $\ell$, the autoregressive policy directly computes contrastive action probability distributions for deriving $\bm{a}_{i}$ (w/ Eq. (2)), whereas the diffusion-based policy requires performing the KDE-PM process on its output predictions to achieve the same goal. Pseudocode for our PCD method is provided in Algorithm 1.
@@
1:A pre-trained robot policy: $\pi_{\theta}(\bm{a}_{i}\mid\bm{o}_{i},\ell)$; initial observation $\bm{o}_{0}$; language command $\ell$; maximum time step $S$.
[…]
@@
4:  Obtain object-masked $\bm{o}_{s}$: $\hat{\bm{o}}_{s}$ $\triangleright$ w/ Track2Mask
5:  Obtain $\pi_{\theta}(\bm{a}_{s}|l,\bm{o}_{s})$ and $\pi_{\theta}(\bm{a}_{s}|l,\hat{\bm{o}}_{s})$ $\triangleright$ w/ KDE-PM for diffusion-based policies
6:  Compute $\pi_{\theta}^{*}(\bm{a}_{s}|l,\bm{o}_{s})$ using Eq. (2)
7:  Sample $\bm{a}_{s}$ from $\pi_{\theta}^{*}(\bm{a}_{s}|l,\bm{o}_{s})$ and execute $\bm{a}_{s}$
8:  $\bm{o}_{s+1}\leftarrow\text{New observation}$ $\triangleright$ Obtain a new observation
@@ 4 Experiments
In this section, we seek to answer the following questions: 1) Can PCD improve robot policies in both simulation and real-world environments?
@@
3) What kinds of spurious correlations can our PCD tackle?
We answer the first question in Sections 4.1 and 4.2, the second question in Section 4.3, and the third question in Section $4.$ Detailed descriptions of the baseline robot policies and the evaluation tasks used in simulation and real-world environments are presented in Appendix C.
Table 1: SIMPLER Performance. Task-specific objects in the initial observation are annotated by artificial Point and Box prompts or the automatic detection results of Grounding DINO (Liu et al., 2024b). The results are the success rate and the 95% confidence interval of 300 trials. As a plug-and-play approach, PCD consistently enhances the three policies by large margins over the 9 tasks.
OpenVLA(Kim et al., 2024) Octo (M. et al., 2024) $\pi_{0}$ (Black et al., 2024)
Tasks Base +PCD Base +PCD Base +PCD
Point Box GDINO Point Box GDINO Point Box GDINO
[…]
+45.2% +36.3% +50.6% +27.5% +26.1% +29.7% +6.6% +7.4% +8.9%
Google
Robot
[…]
@@ 4.1 Simulation Experiments
We perform simulation experiments in SIMPLER (Li et al., 2024), a real-to-sim evaluation environment designed specifically for real-robot policies. SIMPLER accurately reflects real-world performance, enabling reliable assessment of robotic policies. We evaluate 9 tasks across two robot platforms: 5 tasks using the Google Robot and 4 tasks using the WidowX arm.
Experimental Setup. We conduct experiments on top of three diverse robot policies, including the autoregressive policy OpenVLA (Kim et al., 2024), and the diffusion-based policies Octo (base) (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024). We treat these policies as black-boxes, applying PCD solely based on the output action probability distributions for OpenVLA and the output actions for Octo and $\pi_{0}$. For our PCD, we annotate the target object in the initial observation using three manners: artificial Point and Box prompts, and automatic detection results from Grounding DINO (Liu et al., 2024b). The number of the sampled noise vectors NN in KDE-PM is set to 24 for Octo and $\pi_{0}$. We perform ablation studies on the hyperparameter $\alpha$ in Eq. (2), as well as on the object detection strategies and object inpainting paradigms of the devised Track2Mask module in Section 4.3.
Results and Analysis. Table 1 reports the performance of three robot policies (i.e., OpenVLA (Kim et al., 2024), Octo (M. et al., 2024) and $\pi_{0}$ (Black et al., 2024)) integrated w/ or w/o our PCD across a total of 9 tasks from Google Robot and WidowX. From the results in the table, we have the following key observations. 1) Our PCD consistently improves the three baseline policies over the 9 tasks from both the Google Robot and WidowX robotic platforms. On average, PCD boosts the success rates of the autoregressive policy OpenVLA, and the diffusion-based policies Octo and $\pi_{0}$ by up to 50.6%, 29.7% and 8.9%, respectively. 2) Our PCD exhibits stable and outstanding performance across the three object annotation strategies, including artificial Point and Box prompts and automatic object detection with Grounding DINO, underscoring its effectiveness and adaptability. In particular, on the Octo and OpenVLA baseline policies, leveraging the off-the-shelf model Grounding DINO demonstrates better performance than using human Point and Box prompts, while showing comparable results on $\pi_{0}$, highlighting the practical applicability of the PCD method. 3) Compared to OpenVLA and Octo, the baseline model $\pi_{0}$ demonstrates significantly superior performance across all 9 tasks. For instance, OpenVLA and Octo achieve near-zero success rates on “Apple Drawer” task whereas $\pi_{0}$ achieves success rates of 17.0% and 68.7% on these two tasks, respectively. Despite this, from the average results over the 9 tasks, our PCD still enhances the strong baseline $\pi_{0}$ by 8.9%. In summary, the obtained results in the table demonstrate that our PCD approach can be used as a plugin to improve different types of robot policies.
@@ 4.2 Real-world Experiments
Figure 3: Real-world Performance. The target objects in the initial observation are automatically annotated by Grounding DINO (Liu et al., 2024b). PCD delivers a remarkable 108% performance improvement on the baseline, though it incurs a 24% increase in time cost.
@@
Experimental Setup. The simulation experiments show that $\pi_{0}$ significantly outperforms Octo and OpenVLA. Therefore, we employ $\pi_{0}$ as the baseline policy to validate the effectiveness of our PCD approach, as shown in Fig. 3. Considering the generalization limitations of existing robot policies in real-world environments, we first finetune $\pi_{0}$ on downstream real-world tasks before evaluating PCD’s performance. We design 6 manipulation tasks that focus on evaluating the policy’s performance along multiple dimensions: spatial reasoning and interacting with various objects and scenes. The $\pi_{0}$ policy is jointly fine-tuned on the 6 real-world tasks, with each task consisting of 10 training demonstrations. During inference, we conduct 20 trials for each real-world task and randomize the configurations and orientations of task-specific objects for each trial. We use an AGILEX PIPER 6DOF robot arm in our experiments. For our PCD, we leverage the off-the-shelf model Grounding DINO to detect task-specific objects in the initial observation, and use LaMa (Suvorov et al., 2022) to inpaint those objects segmented by the SAM2 (Kirillov et al., 2023) model. All other hyperparameters are kept consistent with the settings applied in the simulation experiments.
@@
Results and Analysis. Fig. 3 reports the success rates as well as time costs of the state-of-the-art robot policy $\pi_{0}$ (Black et al., 2024) integrated w/ or w/o our PCD on 6 real-world tasks. For each task, we record the time cost (in seconds) from the start of the inference process to the successful completion of the task. From the obtained results in the figure, we have several key observations. 1) PCD consistently improves the success rates of the strong baseline policy across six real-world tasks, achieving an average enhancement of 108%. 2) While PCD significantly boosts success rates, it also increases the time cost by 24% on average compared to the baseline policy. However, considering the notable improvement in task completion success rates, this trade-off is acceptable in a wide range of real-world robotic applications. 3) PCD demonstrates consistent performance improvements across tasks of varying complexity. For example, in the challenging "Stack Cube" task, PCD improves the success rate from 0.05 to 0.10, showing its strong adaptability and robustness across diverse scenarios. Additionally, the backgrounds of these six real-world tasks are far more complex than those of the nine simulation tasks, containing numerous distracting elements such as plastic bottles, baskets, and trash cans. Nevertheless, our PCD shows remarkable advantages. It is worth mentioning that we use consistent PCD hyperparameters across tasks in both the simulation and real-world environments, without carefully tuning task-specific PCD hyperparameters (e.g., $\alpha$) for each individual task, leaving room for potential performance improvements through task-specific parameter tuning.
@@ 4.3 Ablation Studies
Effect of Off-the-shelf Object Detection Models. Beyond leveraging artificial Point and Box prompts, the devised Track2Mask module can employ off-the-shelf open vocabulary object detection models to annotate task-specific objects in the initial observation of the trajectory during inference. Fig. 4 (b) presents an ablation study on various automatic object detection models, including the Grounding DINO (Liu et al., 2024b) model discussed in the main paper, as well as the open vocabulary object detection model YOLO World (Cheng et al., 2024) and semantic segmentation model SED (Xie et al., 2024). The output of these models is fed into the SAM2 (Kirillov et al., 2023) model for tracking and segmenting the detected objects across subsequent observations in the trajectory. As can be seen, our PCD consistently improves the baseline policies when integrated with each of the three off-the-shelf models, showcasing its effectiveness and stability. Generally, the Grounding DINO model yields the best results among the three models.
Effect of Object Inpainting Strategies. Once the objects specified in the language instruction of the target task are segmented by the SAM2 model, they are inpainted to create object-masked observations. In this experiment, we investigate the effect of different object inpainting strategies on PCD’s performance. Fig. 4 (c) shows the results of PCD integrated with three inpainting strategies: Telea (Telea, 2004) and Navier-Stokes (Bertalmio et al., 2001) and LaMa (Suvorov et al., 2022). As indicated in the figure, our PCD approach consistently improves the baseline policies across all strategies, showcasing its effectiveness and stability. Furthermore, PCD demonstrates low sensitivity to variations in the three object inpainting strategies, further validating its applicability and robustness. Notably, LaMa delivers the best results across all three baseline policies.
@@ 4.4 Robustness to Various Kinds of Spurious Correlations
The central concept of our proposed PCD approach in this work is to overcome the adverse effects of spurious features on the decision-making of robot policies. So, what types of spurious correlations can PCD effectively address? To answer the question, we evaluate the robustness of the OpenVLA (Kim et al., 2024) and $\pi_{0}$ (Black et al., 2024) policies to task scene variations in both simulation and real-world environments. As illustrated in Fig. 5, we report the performance of the two baseline policies integrated w/ or w/o our proposed PCD approach in different testing scenarios with unseen spatial relationships, brightness, distractors and table textures within visual observations. The results shown in the figure lead to the following observations. 1) Consistent with the previous experimental results, our PCD approach significantly improves the performance of the baseline policies on both simulation and real-world tasks (Orange line vs. Blue line). 2) Both baseline policies experience substantial drops in performance when testing scenarios are modified. For example, altering the brightness of visual observations results in 48% and 75% decreases in success rates for the two policies, respectively. 3) PCD consistently alleviates the performance degradation of the baseline policies across all unseen testing scenarios. Remarkably, after integrating PCD, the two baseline policies even performs better in 4/10 of unseen scenarios than in the original training scenarios. The results in the figure reveal that our proposed PCD approach can serve as a plugin to mitigate the adverse effects of various types of spurious correlations on the generalization of robot policies.
Figure 5: Performance of baseline policies integrated w/ or w/o our proposed PCD approach in unseen testing scenarios. Dotted lines represent the original performance. OpenVLA (Kim et al., 2024) and $\pi_{0}$ (Black et al., 2024) are used as baselines in simulation and real-world environments, respectively. PCD effectively mitigates the side effects of various types of spurious correlations, boosting the robot policy’s generalization to novel scenarios.